OpenStudy (anonymous):

A dentist's drill starts from rest. After 2.90 s of constant angular acceleration, it turns at a rate of 2.49 x 10^4 rev/min. (a) Find the drill's angular acceleration. rad/s2 (along the axis of rotation) (b) Determine the angle through which the drill rotates during this period. I have the answer I just don't know how to get it. Fan and Medal if you can explain the rotational motion equations

OpenStudy (michele_laino):

The angular acceleration is, by definition, the ratio between the variation of angular velocity and the time within that variation of angular velocity is occurred

OpenStudy (michele_laino):

the equation of motion is: \[\frac{ d ^{2}\theta }{dt ^{2} }=\alpha \] where \[\alpha \] is your angular acceleration